OpenStudy (anonymous):

Which of the following is a polynomial with roots –2, –2i, and 2i? A. x^3 – 4x^2 + 4x – 16 B. x^3 + 4x^2 + 4x + 16 C. x^3 + 2x^2 + 4x + 8 D. x^3 – 2x^2 + 4x – 8

OpenStudy (anonymous):

your job is to multiply \[(x+2)(x+2i)(x-2i)\]

OpenStudy (anonymous):

second part is easy, because \[(x+bi)(x-bi)=x^2+b^2\]

OpenStudy (anonymous):

so your real job is to multiply \[(x+2)(x^2+4)\] and see what you get

OpenStudy (anonymous):

wait, how do you multiply x and -2i?

OpenStudy (anonymous):

x^3 + 4x + 2x^2 + 8 x^3 + 2x^2 + 4x + 8

OpenStudy (anonymous):

\[(x+2i)(x-2i)=x^2+2ix-2ix-2i\times 2i=x^2-4i^2=x^2+4\]

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (anonymous):

yes you have it yw